OpenStudy (tywower):

Line L has equation 2x - 3y = 5. Line M passes through the point (2, -10) and is perpendicular to line L. Determine the equation for line M. A) 2x + 3y = -14 B) 2x - 3y = -26 C) 3x + 2y = -14 Eliminate D) 3x - 2y = -26

OpenStudy (tywower):

@Nnesha

OpenStudy (tywower):

@Nnesha

OpenStudy (tywower):

@Nnesha

OpenStudy (tywower):

@Nnesha

OpenStudy (tywower):

@Nnesha

OpenStudy (tywower):

@Nnesha

OpenStudy (tywower):

@perl

OpenStudy (tywower):

@naledimunetsi

OpenStudy (tywower):

@quickstudent

OpenStudy (tywower):

@quickthing

OpenStudy (tywower):

@GeniousCreation

OpenStudy (quickthing):

Sorry, I am terrible at math.

OpenStudy (tywower):

@Nnesha

OpenStudy (tywower):

@Nnesha

OpenStudy (tywower):

@Nnesha

OpenStudy (tywower):

@GeniousCreation

OpenStudy (tywower):

YAY your finally here

OpenStudy (tywower):

pls

Nnesha (nnesha):

what's the slope of this line 2x -3y = 5 ???

OpenStudy (tywower):

idk 2

Nnesha (nnesha):

y = mx + b slope intercept form where m is slope 2x -3y = 5 solve this equation for y

OpenStudy (tywower):

@GeniousCreation

OpenStudy (tywower):

omg please walk me through it

Nnesha (nnesha):

2x - 3y = 5 subtract 2x both side

OpenStudy (tywower):

ok

OpenStudy (tywower):

and then

Nnesha (nnesha):

let me know what you get after subtracting 2x both side ?

OpenStudy (tywower):

-3y=5-2x

Nnesha (nnesha):

now left side is *-3 times y* do opposite of multiply to cancel out -3

OpenStudy (tywower):

idk please can u not

OpenStudy (anonymous):

Is the 3 a negative number?

OpenStudy (tywower):

yes

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @tywower idk please can u not \(\color{blue}{\text{End of Quote}}\) what ?? what is opposite of multiply ?

OpenStudy (anonymous):

Divide!

OpenStudy (tywower):

division

Nnesha (nnesha):

2nd way to find equation substitute (2 , -5) in all answer choices if you get both equal sides equal of equation then that equation would be your answer

Nnesha (nnesha):

yes so divide both side by -3

Nnesha (nnesha):

 what;s the slope ?